If the perimeter of a rectangle is 20 ft and the width is 6ft, find its length.[hint the perimeter of a rectangle is given by P=2L+2W.]
Answer by checkley71(8403)

20=2L+2*6
20=2L+12
2L=20-12
2L=8
L=8/2
L=4 ANSWER FOR THE LENGTH.
PROOF
20=2*4+12
20=8+12
20=20